RAM , six years old, is twice as old as his brother. How old will RAM be when he is one and a half times as old as his brother?
Answer by checkley79(3341)

RAM=6
BROTHER=6/2=3
6+X=1.5(3+X)
6+X=4.5+1.5X
X-1.5X=4.5-6
-.5X=-1.5
X=-1.5/-.5
X=3 YEARS FROM NOW RAM WILL BE 1.5 TIMES HIS BROTHERS AGE.
pROOF:
6+3=1.5(3+3)
9=1.5*6
9=9